OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (2, 7) and (-6, -1)?

OpenStudy (anonymous):

1/2(2+(-6),7+(-1))=1/2(-4,6)=(-2,3)

OpenStudy (anonymous):

great, thanks! :)

OpenStudy (anonymous):

yw